# Improved treatment of relativistic effects in linear augmented plane wave

### v1

@@ Acknowledgements.
In the spherically symmetric potential the $j=\ell-1/2$ and $j=\ell+1/2$ levels are degenerate. If both levels are completely filled, which holds e.g. for the fully occupied $6p_{1/2}$ and $6p_{3/2}$ semicore states of the actinides, then, according to the generalized Unsöld theorem, the sum of squares of all angular components of the $j=\ell-1/2$ and $j=\ell+1/2$ state is an invariant independent of the polar angles. In this case the partial weights of the $j=\ell-1/2$ and $j=\ell+1/2$ levels in the total electron density are $w_{\ell,1}=\ell/(2\ell+1)$ and $w_{\ell,2}=(\ell+1)/(2\ell+1)$, correspondingly. Therefore, these weight factors can be used to correct the electron density, constructed from the averaged large component, by adding the extra terms associated with the small components $Q_{-\ell}(r)$ and $Q_{\ell+1}$ of these states, i.e.
$\rho_{\ell}(r)=\rho_{\ell,0}\left[P_{\ell}^{2}(r)+w_{\ell,1}Q_{-\ell}^{2}(r)+w_{\ell,2}Q_{\ell+1}^{2}(r)\right],$ (22)
where $\rho_{\ell,0}$ is the $\ell$-partial charge obtained for the averaged large radial component $P_{\ell}$. In LAPW method there are two radial functions – $P_{\ell}$ and $\dot{P}_{\ell}=dP_{\ell}/dE$. Therefore, in addition to $\rho_{\ell}(r)$, Eq. (22) we also take into account the electron density $\rho^{E}_{\ell}(r)$, associated with the energy derivatives $\dot{Q}_{-\ell}(r)$ and $\dot{Q}_{{\ell}+1}$ of the small components $Q_{-\ell}(r)$ and $Q_{{\ell}+1}$, i.e.
$\rho_{\ell,\,tot}(r)=\rho_{\ell}(r)+\rho^{E}_{\ell}(r),$ (23)
where
$\rho^{E}_{\ell}(r)=\rho^{E}_{\ell,0}\left[\dot{P}_{\ell}^{2}(r)+w_{\ell,1}\dot{Q}_{-\ell}^{2}(r)+w_{\ell,2}\dot{Q}_{\ell+1}^{2}(r)\right].\quad$ (24)
Here $\rho^{E}_{\ell,0}$ is the partial charge for $\dot{P}_{\ell}$. The electron densities (23) are involved in the calculation variants denoted as avD(Q) in tables 69. The total valence values of $\rho(r)$, averaged over the actinide nuclei of finite size, are denoted in the last column of these Tables as $\rho_{nuc}$. They can be compared with the electron densities obtained only with the large components $P_{\ell}$ and $\dot{P}_{\ell}$. Practically all increase (96%) of $\rho_{nuc}$ is due to the small components of the semicore $6p$ states.
@@
In our study we tend to use relatively large values of $R_{MT}$, which is related to our choice to work with a large number of radial points inside MT-sphere (4000-4200 points for actinides). The large number of radial points improves the accuracy of LAPW basis functions avoiding possible linear dependence. On the other hand, the large number of radial points improves the description of 80 core electrons for actinides, which is quite a big quantity requiring very accurate radial treatment.

### v2

@@ Acknowledgements.
In the spherically symmetric potential the $j=\ell-1/2$ and $j=\ell+1/2$ levels are degenerate. If both levels are completely filled, which holds e.g. for the fully occupied $6p_{1/2}$ and $6p_{3/2}$ semicore states of the actinides, then, according to the generalized Unsöld theorem, the sum of squares of all angular components of the $j=\ell-1/2$ and $j=\ell+1/2$ state is an invariant independent of the polar angles. In this case the partial weights of the $j=\ell-1/2$ and $j=\ell+1/2$ levels in the total electron density are $w_{\ell,1}=\ell/(2\ell+1)$ and $w_{\ell,2}=(\ell+1)/(2\ell+1)$, correspondingly. Therefore, these fixed weight factors can be used to correct the electron density, constructed from the averaged large component, by adding the extra terms associated with the small components $Q_{-\ell}(r)$ and $Q_{\ell+1}$ of these states. As a simple example, we first consider the spherically symmetric charge density $\rho_{\ell,0}P_{\ell}^{2}(r)$ obtained for the averaged large radial component $P_{\ell}$. Here $\rho_{\ell,0}$ is the computed $\ell$-partial charge, associated with $P_{\ell}$. In this case the full electron density $\rho_{\ell}(r)$, including two small components, is
$\rho_{\ell}(r)=\rho_{\ell,0}\left[P_{\ell}^{2}(r)+w_{\ell,1}Q_{-\ell}^{2}(r)+w_{\ell,2}Q_{\ell+1}^{2}(r)\right].\quad$ (22)
In LAPW method there are two radial functions: $P_{\ell}$ and $\dot{P}_{\ell}=dP_{\ell}/dE$. Therefore, in addition to $\rho_{\ell}(r)$, Eq. (22) we have to take into account the electron density $\rho^{E,E}_{\ell}(r)$, associated with the squares of the energy derivatives $\dot{Q}_{-\ell}(r)$ and $\dot{Q}_{{\ell}+1}$ of the small components $Q_{-\ell}(r)$ and $Q_{{\ell}+1}$. There is also a mixed contribution $\rho^{E}_{\ell}(r)$, which includes terms like $Q_{-\ell}(r)\dot{Q}_{-\ell}(r)$ and $Q_{\ell+1}(r)\dot{Q}_{\ell+1}(r)$. The full electron density then reads as
$\rho_{\ell,\,tot}(r)=\rho_{\ell}(r)+\rho^{E}_{\ell}(r)+\rho^{E,E}_{\ell}(r).$ (23)
Here
$\rho^{E,E}_{\ell}(r)=\rho^{E,E}_{\ell,0}\left[\dot{P}_{\ell}^{2}(r)+w_{\ell,1}\dot{Q}_{-\ell}^{2}(r)+w_{\ell,2}\dot{Q}_{\ell+1}^{2}(r)\right],$
where $\rho^{E,E}_{\ell,0}$ is the partial charge for $\dot{P}_{\ell}$, and we have omitted the analogous expression for $\rho^{E}_{\ell}$. In tables 69 the calculation variants, involving the full electron densities (23 with small components, are denoted as avD(Q). The full valence electron densities $\rho(r)$, averaged over the actinide nuclei of finite size in these variants, are given in the last columns of these Tables as $\rho_{nuc}$. They can be compared with the electron densities obtained only with the large components $P_{\ell}$ and $\dot{P}_{\ell}$ in other calculations. Practically all increase (96%) of $\rho_{nuc}$ is due to the small components of the semicore $6p$ states.
@@
In our study we tend to use relatively large values of $R_{MT}$, which is related to our choice to work with a large number of radial points inside MT-sphere (4000-4200 points for actinides). The large number of radial points increases the accuracy of LAPW basis functions avoiding possible linear dependence. On the other hand, a large number of radial points improves the description of the 80 core electrons for actinides with high electron density in the nucleus, which decreases rapidly with radius, requiring a very accurate radial representation.